Together John and Michael can paint a wall in 22 minutes. Alone John needs 15 minutes longer to paint the wall than Michael needs. How much time does John need to paint the wall by himself?
Answer by richwmiller(17219)

See
http://www.algebra.com/algebra/homework/Rate-of-work-word-problems/Rate-of-work-word-problems.faq.question.910120.html
1/x+1/(x+15)=1/22
(22*x*(x+15))/x+(22*x*(x+15))/(x+15)=(22*x*(x+15))/22
(22*(x+15))+22*x=x*(x+15)
22*x+330+22*x=x^2+15*x
44*x+330=x^2+15*x
x^2-29*x-330=0
Positive useful answer:
x=14.5+23.2432786
One answer is x=37.7
x+15=15+37.7
x+15=52.7